solve the system of equations

3x−4y=8
​18x−5y=10
​​

Answer by Boreal(15235)

3x−4y=8
​18x−5y=10
eliminate x by multiplying the top equation by -6 then adding the two.
-18x+24y=-48
18x-5y=10
19y=-38
y=-2
Substitute into first equation
3x-4(-2)=8
3x+8=8
x=0
Check (0.-2) in the second equation.
0-5(-2)=10. It checks.